HETAV PATEL

Grade 11,

h.c. verma question- A railway track( laid of iron) is laid in winter when temp. is 18 C.the track consists of sections of 12m placed one after the other.How much gap shud be kept b/w two such sections so that there is no compression in summer at temp. 48 C ?coeff. of linear expansion of iron=1.1x10(raise to -6). here i think the gap shud be twice of linear expansion as both the rods kept in contact will expand but the answer given is equal to linear expansion. please someone reply soon...

h.c. verma question-
A railway track( laid of iron) is laid in winter when temp. is 18 C.the track consists of sections of 12m placed one after the other.How much gap shud be kept b/w two such sections so that there is no compression in summer at temp. 48 C ?coeff. of linear expansion of iron=1.1x10(raise to -6). here i think the gap shud be twice of linear expansion as both the rods kept in contact will expand but the answer given is equal to linear expansion. please someone reply soon...

Grade:12th Pass

1 Answers

Chetan Mandayam Nayakar
12 years ago

let 'x' be the linear expansion. linear expansion takes place on BOTH SIDES of each section. let 'a' be the gap on both sides. Linear expansion is x/2 into the gap on eaxh side. In addition adjacent section on each side expands into the gap by x/2,so gap=x/2 +x/2 =x=linear expansion.